OpenStudy (anonymous):

Let f(x) = 4x2 + x + 1 and g(x) = x2 – 2. Find g(f(x)). Show each step of your work.

OpenStudy (anonymous):

@Baby_Bear69

OpenStudy (anonymous):

is my work right? F(x)=4x^2+x+1 G(x)=x^2-2 G(f(x))=(4x^2+X+1)^2-2 G(f(x))=16X^4+X^2+1^2+2.4X^2+X+2.4X^2+1+2X-2 G(f(x))= 16X^4 +8X^3+9x^2+2X-1

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

\[\large \bf gof(x)=g(f(x))\] Simply put function f(x) into function g(x) in place of variable.

OpenStudy (anonymous):

so 4x^2+x+1=x^2-2

OpenStudy (mayankdevnani):

nope Lemme tell you

OpenStudy (mayankdevnani):

\[\large \bf f(x)=4x^2+x+1\] \[\large \bf g(x)=x^2-2\] \[\large \bf g(f(x))=(4x^2+x+1)^2-2\] in additional , \[\large \bf f(g(x))=4(x^2-2)^2+(x^2-2)+1\]

OpenStudy (mayankdevnani):

hope you understand ! @Motown117

OpenStudy (anonymous):

ok i see where i went wrong

OpenStudy (mayankdevnani):

:D

OpenStudy (anonymous):

thank you for the clarification and the help

OpenStudy (mayankdevnani):

your welcome :)